# Exact Bayesian Inference for Markov Switching Diffusions

### v1

@@ 3 Exact Posterior Sampling using Barker-within-Gibbs
$(V_{R},Z,Y):\quad$ $\pi(v_{r},z,y|v_{s},\theta,\lambda)\propto\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)\pi(y_{\ddot{\tau}}|y_{\dot{\tau}},\lambda),$ (14)
where, keeping in mind that the set of jumps $R$ and event times $T$ follow deterministically from $Y$, and that $R$ is almost surely finite, the dominating measure of the second full conditional is the product measure $\mathbb{L}(\mathop{}\!\mathrm{d}y)\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\mathbb{B}_{(\dot{\tau},\ddot{\tau})}(\mathop{}\!\mathrm{d}z_{(\dot{\tau},\ddot{\tau})})\prod_{\dot{r}\in r}\mathrm{Leb}(\mathop{}\!\mathrm{d}v_{\dot{r}})$, which is invariant in $\theta$. This blocking offers various opportunities for exploiting conditional independence and tuning proposals. Indeed, the $(\Theta,\Lambda)$-update decomposes into the independent updates
@@
The $\Lambda$-update is conjugate for the class of priors discussed in Section 3.4, and may thus be sampled exactly. In addition, if the SDE depends on a separate set of parameters for each latent mode, the updates to those parameters may also be carried out independently. We exploit that structure in the models seen in Sections 5 and 6. Having specified the full conditional distributions, we proceed with a presentation of the Bernoulli MCMC approach in Section 3.1 before presenting algorithms that update the respective full conditionals in Sections 3.2 and 3.3.
@@ 3.1 Retrospective Simulation for Diffusion Inference
In what follows, both updates of the Gibbs sampler are carried out by accept-reject coin flips, where the acceptance probability depends on the intractable complete likelihood $\pi(v_{\tau\mathbin{\backslash}\{0\}},z|v_{0},y,\theta)$. As originally suggested in the Ito diffusion context by gonccalves2017barker and later applied to jump diffusions in gonccalves2023exact, such evaluations can be entirely avoided without affecting the dynamics of the Markov chain. This is accomplished through the conjunction of three insights: Firstly, we may construct events with probabilities proportional to $\pi(v_{\tau\mathbin{\backslash}\{0\}},z|v_{0},y,\theta)$ that depend only on a finite number of evaluations of $z$. In the instance of diffusions, this is known as the Poisson coin algorithm (beskos2006retrospective). Secondly, $z$ may be left indeterminate until specific evaluations are needed to propagate the Markov chain. This skeleton of $z$ is then simulated retrospectively at the required locations. Finally, using Bernoulli factory techniques, we may construct an accept-reject coin flip from Poisson coin flips.
Before delving deeper into the Bernoulli MCMC approach, we briefly recall some fundamentals of accept-reject MCMC. Suppose we are targeting the density $\pi(a)$ with $A$ taking values in the state space $\mathcal{A}$ and a proposal density $\kappa(a^{\dagger}|a)$. The Metropolis-Hastings (M-H) acceptance probability for a proposal $a^{\dagger}$ is $\alpha_{\text{MH}}(a,a^{\dagger})=1\land\{\pi(a^{\dagger})\kappa(a|a^{\dagger})\}/\{\pi(a)\kappa(a^{\dagger}|a)\}$, and its repeated application results in a Markov chain with stationary density $\pi(a)$ under mild conditions on $\kappa$ due to the detailed balance property. It may also be used to update a full conditional distribution within a Gibbs sampler when direct sampling thereof is not possible, which is known as Metropolis-within-Gibbs. There are other acceptance probabilities that preserve detailed balance, such as accepting with odds $\alpha_{\text{B}}(a,a^{\dagger})/\alpha_{\text{B}}(a^{\dagger},a)=\{\pi(a^{\dagger})\kappa(a|a^{\dagger})\}/\{\pi(a)\kappa(a^{\dagger}|a)\}$, as proposed by barker1965monte. These are rarely applied since M-H induces smaller autocorrelation in the Markov chain than other reversible acceptance probabilities, as shown in peskun1973optimum. Nonetheless, Barker’s acceptance probability takes on a special role in the intractable likelihood context. Firstly, as shown in latuszynski2013clts, when estimating a test function $f(a)$ by its ergodic average under each Markov chain, the asymptotic variances of the two estimators are within a factor of 2, so any problem that may be addressed by M-H is also solvable with Barker. Secondly, the Barker accept-reject coin is more amenable to being constructed through a Bernoulli factory. To expand on the latter, suppose that $\alpha_{\text{B}}$ admits the factorization
@@
where $c_{1},c_{2}$ are tractable constants and $p_{1},p_{2}\in[0,1]$ are potentially intractable, but for which we can simulate coins with equivalent probability, e.g. by way of an unbiased estimator. Then, provided a stream of coin flips of heads-probability $p_{1}$ and another of heads-probability $p_{2}$, the 2-coin algorithm generates coins of probability $\alpha_{\text{B}}(a,a^{\dagger})$ (gonccalves2017barker). We summarize its operation in Figure 2. Then, on a high level, the Bernoulli MCMC algorithm generates $p_{1}$- and $p_{2}$-coins through the Poisson coin method, which we summarize in Supplement D, revealing the diffusion path retrospectively to the required resolution for the coin flips. On that basis, the $\alpha_{\text{B}}$-coins are simulated by the 2-coin algorithm, without ever evaluating $\alpha_{\text{B}}$. Indeed, even in the conventional form, the evaluation of $\alpha_{\text{B}}$ is entirely ancillary, since it is merely used to determine the binary event $\{\alpha_{\text{B}}(a,a^{\dagger})>U\}$ for $U\sim\operatorname{\mathrm{Uniform}}\left[0,1\right]$. How the accept-reject coin is constructed does not affect the dynamics of the Markov chain. On the other hand, the runtime of the 2-coin algorithm is highly dependent on the chosen factorization. Indeed, the number of loops in the 2-coin algorithm is a geometric random variable with expectation
@@
which diverges as $p_{1},p_{2}\to 0$. This is particularly relevant to the $\theta$-update described in Section 3.3, and extensions to the standard 2-coin approach are required to prevent the computational cost from diverging as $\omega$ increases. Of particular relevance is the divide-and-conquer 2-coin algorithm, proposed by stumpf2024bf and applied in Section 3.3, and the Portkey 2-coin algorithm, proposed by vats2020efficient and applied to the experiments in Sections 5 and 6.
@@
Figure 2: Probability flow diagram of the vanilla 2-coin algorithm. Nodes ($F_{0}$, $F_{1}$, $F_{2}$) refer to coin flips, edges give the probabilities of moving to the corresponding node.
@@
In order to simulate $p_{1}$- and $p_{2}$-coins in the broadest set of Markov switching models, the skeleton $\breve{z}_{(\dot{\tau},\ddot{\tau})}$ of the bridge $z_{(\dot{\tau},\ddot{\tau})}$ must satisfy three requirements. Firstly, it must incorporate all previously evaluated locations on the path. Secondly, it must provide lower and upper bounds
$-\infty<{z}^{\downarrow}_{(\dot{\tau},\ddot{\tau})}\leq z_{t}\leq{z}^{\uparrow}_{(\dot{\tau},\ddot{\tau})}<\infty,\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (19)
@@
which can then be propagated to bounds on $\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\varphi_{\theta}(\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),y_{\dot{\tau}})$. Thirdly, it must be possible to recursively simulate $z_{(\dot{\tau},\ddot{\tau})}$ at an additional finite set of times according to the initial proposal distribution, conditional on the skeleton $\breve{z}_{(\dot{\tau},\ddot{\tau})}$. In this paper, the proposal distribution will always be the dominating measure $\mathbb{B}_{(\dot{\tau},\ddot{\tau})}$. Modulo these requirements, any method may in principle be used to generate the coins. For instance, in the $\epsilon$-strong simulation framework of pollock2016exact, the skeleton representation is
$\breve{z}_{(\dot{\tau},\ddot{\tau})}=\{z_{t}:t\in\mathcal{S}_{(\dot{\tau},\ddot{\tau})}\}\cup\{({\check{z}}^{\downarrow}_{(\dot{t},\ddot{t})},{\check{z}}^{\uparrow}_{(\dot{t},\ddot{t})},{\hat{z}}^{\downarrow}_{(\dot{t},\ddot{t})},{\hat{z}}^{\uparrow}_{(\dot{t},\ddot{t})}):(\dot{t}\sim\ddot{t})\in\mathcal{S}_{(\dot{\tau},\ddot{\tau})}\}$ (20)
where $\mathcal{S}_{(\dot{\tau},\ddot{\tau})}$ is the set of times at which $z_{(\dot{\tau},\ddot{\tau})}$ has been previously evaluated, $\check{z}_{(\dot{t},\ddot{t})}=\inf_{t\in(\dot{t},\ddot{t})}z_{t}$, $\hat{z}_{(\dot{t},\ddot{t})}=\sup_{t\in(\dot{t},\ddot{t})}z_{t}$, and $({\check{z}}^{\downarrow}_{(\dot{t},\ddot{t})},{\check{z}}^{\uparrow}_{(\dot{t},\ddot{t})})$ are bounds such that $\check{z}_{(\dot{t},\ddot{t})}\in[{\check{z}}^{\downarrow}_{(\dot{t},\ddot{t})},{\check{z}}^{\uparrow}_{(\dot{t},\ddot{t})}]$ with probability 1. Then, $z_{(\dot{\tau},\ddot{\tau})}$ is trivially bounded within $(\min_{(\dot{t}\sim\ddot{t})\in\mathcal{S}_{(\dot{\tau},\ddot{\tau})}}{\check{z}}^{\downarrow}_{(\dot{t},\ddot{t})},\max_{(\dot{t}\sim\ddot{t})\in\mathcal{S}_{(\dot{\tau},\ddot{\tau})}}{\hat{z}}^{\uparrow}_{(\dot{t},\ddot{t})})$. Another representation was devised for the EA3 algorithm in beskos2008factorisation. In what follows, we merely presume that some lower and upper bounds ${z}^{\downarrow}_{(\dot{\tau},\ddot{\tau})}$ and ${z}^{\uparrow}_{(\dot{\tau},\ddot{\tau})}$ are available. Observing that $x_{t}\in\mathcal{I}_{(\dot{\tau},\ddot{\tau})}=[\rho_{\theta}(y_{\dot{\tau}}){z}^{\downarrow}_{(\dot{\tau},\ddot{\tau})}+\eta_{\theta}(v_{\dot{\tau}})\land\eta_{\theta}(v_{\ddot{\tau}}),\rho_{\theta}(y_{\dot{\tau}}){z}^{\uparrow}_{(\dot{\tau},\ddot{\tau})}+\eta_{\theta}(v_{\dot{\tau}})\lor\eta_{\theta}(v_{\ddot{\tau}})]$, we define
@@
${\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\inf_{a\in\mathcal{I}_{(\dot{\tau},\ddot{\tau})}}\varphi_{\theta}(a,y_{\dot{\tau}}),\quad{\tilde{\varphi}}^{\uparrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\sup_{a\in\mathcal{I}_{(\dot{\tau},\ddot{\tau})}}\varphi_{\theta}(a,y_{\dot{\tau}}),$ (21)
@@
such that ${\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\leq\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\leq{\tilde{\varphi}}^{\uparrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$. Solving for the infimum and supremum, or bounds thereon, is a model-specific task, but may be accomplished automatically by symbolic algebra packages.
@@ 3.2 Hidden Data Update
where $Z_{(\dot{\tau},\ddot{\tau})}^{\dagger}\sim\mathbb{B}_{(\dot{\tau},\ddot{\tau})}$ and $\kappa(z_{(\dot{\tau},\ddot{\tau})}^{\dagger})=1$, with respect to $\mathbb{B}_{(\dot{\tau},\ddot{\tau})}$. Using a Brownian bridge proposal ensures that the path skeletons can be simulated and extended, e.g. by the EA algorithm. $Y^{\dagger}$ is proposed independently from its prior distribution, i.e. $\kappa(y^{\dagger})=\pi(y^{\dagger}|\lambda)$. This is simply the forward measure of the jump process, and easily simulated from as in hobolth2009simulation. Given $Y^{\dagger}$, the proposal for $V_{r^{\dagger}}^{\dagger}$ is most readily understood in terms of $X_{r^{\dagger}}^{\dagger}=\eta_{\theta}(V_{r^{\dagger}}^{\dagger})$. We propose $X_{r^{\dagger}}^{\dagger}$ according to the dominating SDE $\mathop{}\!\mathrm{d}X_{t}=\rho_{\theta}(Y_{t}^{\dagger})\mathop{}\!\mathrm{d}W_{t}$ with induced conditional measure $\mathbb{M}|(x_{s},y^{\dagger},\theta)$. By the Markov property,
@@
$(t,x_{t})\mapsto\left\lparen\int_{\dot{s}}^{t}\rho_{\theta}^{2}(y_{u})\mathop{}\!\mathrm{d}u,x_{t}\right\rparen$ ) (24)
@@
Thus, the proposal density on $V_{r^{\dagger}}^{\dagger}$ is given by $\kappa(v_{r^{\dagger}}^{\dagger}|v_{s},y^{\dagger})=\prod_{(\dot{s}\sim\ddot{s})\in s}\kappa(v_{r^{\dagger}\cap(\dot{s},\ddot{s})}^{\dagger}|v_{\{\dot{s},\ddot{s}\}},y_{[\dot{s},\ddot{s}]}^{\dagger})$. We give the step-by-step routine below.
Algorithm 1 Algorithm for generating proposal from $\kappa(v_{r\cap(\dot{s},\ddot{s})}|v_{\{\dot{s},\ddot{s}\}},y_{[\dot{s},\ddot{s}]})$. $\mathbb{W}$ denotes the Wiener mesure.
@@
$d_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})e^{(\dot{\tau}-\ddot{\tau}){\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})},$ (27)
@@
$q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\exp\left\{\int_{\dot{\tau}}^{\ddot{\tau}}{\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\mathop{}\!\mathrm{d}t\right\},$ (28)
@@
such that $\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta^{\dagger})=d_{\theta}(z_{(\dot{\tau},\ddot{\tau})},v_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}})q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$, and where $q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\in[0,1]$. On that basis, the acceptance odds can be expressed as

### v2

@@ 3 Exact Posterior Sampling using Barker-within-Gibbs
$(V_{R},Z,Y):\quad$ $\pi(v_{r},z,y|v_{s},\theta,\lambda)\propto\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)\pi(y_{\ddot{\tau}}|y_{\dot{\tau}},\lambda).$ (14)
Keeping in mind that the set of jumps $R$ and event times $T$ follow deterministically from $Y$, and that $R$ is almost surely finite, the dominating measure of the second full conditional is the product measure $\mathbb{L}(\mathop{}\!\mathrm{d}y)\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\mathbb{B}_{(\dot{\tau},\ddot{\tau})}(\mathop{}\!\mathrm{d}z_{(\dot{\tau},\ddot{\tau})})\prod_{\dot{r}\in r}\mathrm{Leb}(\mathop{}\!\mathrm{d}v_{\dot{r}})$, which is invariant in $\theta$. This blocking offers various opportunities for exploiting conditional independence and tuning proposals. In fact, the $(\Theta,\Lambda)$-update decomposes into the independent updates
@@
Further simplifications arise if the SDE depends on a separate set of parameters for each latent mode. In that case, the updates to those parameters may also be carried out independently. We exploit that structure in the models seen in Sections 5 and 6.
Each of the Gibbs updates poses distinct challenges. For the $(V_{R},Z,Y)$-update, the construction and simulation of the proposal is fairly intricate, and the accept-reject procedure is a Bernoulli factory that implements Barker’s algorithm within Gibbs [3], which just as Metropolis-within-Gibbs leaves the full conditional distribution invariant. We introduce the Bernoulli factory approach to diffusion inference in Section 3.1 before specifying the proposal and the factory for the $(V_{R},Z,Y)$-update in Section 3.1. The $\Theta$-update again relies on a Barker-within-Gibbs update with a simple random walk proposal, though specific scalability challenges arise when using Bernoulli factories in such an instance. We describe and address those challenges in Section 3.3. Finally, the $\Lambda$-update is conjugate for the class of priors discussed in Section 3.4, and may thus be sampled exactly from the full conditional. We close the Section with a discussion of various practical implementation challenges in Section 4.3.
@@ 3.1 Retrospective Simulation for Diffusion Inference
[…]
Before delving deeper into the Bernoulli MCMC approach, we briefly recall some fundamentals of accept-reject MCMC. Suppose we are targeting the density $\pi(a)$ with $A$ taking values in the state space $\mathcal{A}$ and a proposal density $\kappa(a^{\dagger}|a)$. The Metropolis-Hastings (MH) acceptance probability for a proposal $a^{\dagger}$ is $\alpha_{\text{MH}}(a,a^{\dagger})=1\land\{\pi(a^{\dagger})\kappa(a|a^{\dagger})/\{\pi(a)\kappa(a^{\dagger}|a)\}\}$, and its repeated application to random proposals results in a Markov chain with stationary density $\pi(a)$ under mild conditions on $\kappa$ due to the detailed balance or reversibility property. It may also be used to update a full conditional distribution within a Gibbs sampler when direct sampling thereof is not possible, which is known as Metropolis-within-Gibbs. There are other acceptance probabilities that preserve detailed balance, such as accepting with odds $\alpha_{\text{B}}(a,a^{\dagger})/\alpha_{\text{B}}(a^{\dagger},a)=\pi(a^{\dagger})\kappa(a|a^{\dagger})/\{\pi(a)\kappa(a^{\dagger}|a)\}$, as proposed by [3]. These are rarely applied since MH induces smaller autocorrelation in the Markov chain than other reversible acceptance probabilities, as shown in [38]. Nonetheless, Barker’s acceptance probability takes on a special role in the intractable likelihood context. Firstly, as shown in [30], when estimating a test function $f(a)$ by its ergodic average under each Markov chain, the asymptotic variances of the two estimators are within a factor of 2, so any problem that may be addressed by MH is also solvable with Barker. Secondly, the Barker accept-reject coin is more amenable to being constructed through a Bernoulli factory. To expand on the latter, suppose that $\alpha_{\text{B}}$ admits the factorization
@@
where $c_{1},c_{2}$ are tractable constants and $p_{1},p_{2}\in[0,1]$ are potentially intractable, but for which we can simulate coins with equivalent probability, e.g. by way of an unbiased estimator. Then, provided a stream of coin flips of heads-probability $p_{1}$ and another of heads-probability $p_{2}$, the 2-coin algorithm generates coins of probability $\alpha_{\text{B}}(a,a^{\dagger})$ [21]. We summarize its operation in Figure 2. Then, on a high level, the Bernoulli MCMC algorithm generates $p_{1}$- and $p_{2}$-coins through the Poisson coin method, which we summarize in Supplement D, revealing the diffusion path retrospectively to the resolution required for the coin flips. On that basis, the $\alpha_{\text{B}}$-coins are simulated by the 2-coin algorithm, without ever evaluating $\alpha_{\text{B}}$. In fact, the evaluation of $\alpha_{\text{B}}$ is entirely ancillary, since it is merely used to determine the binary event $\{\alpha_{\text{B}}(a,a^{\dagger})>U\}$ for $U\sim\operatorname{\mathrm{Uniform}}\left[0,1\right]$. How the accept-reject coin is constructed does not affect the dynamics of the Markov chain. On the other hand, the runtime of the 2-coin algorithm is highly dependent on the chosen factorization: the number of loops in the 2-coin algorithm is a geometric random variable with expectation
@@
which diverges as $p_{1},p_{2}\to 0$. This is critical to the $\theta$-update described in Section 3.3, and extensions to the standard 2-coin approach are required to prevent the computational cost from diverging exponentially as $\omega$ increases. Of particular relevance is the divide-and-conquer 2-coin algorithm, proposed by [44] and applied in Section 3.3, and the Portkey 2-coin algorithm, proposed by [45] and applied to the experiments in Sections 5 and 6.
@@
Figure 2: Probability flow diagram of the vanilla 2-coin algorithm. Nodes ($C_{0}$, $C_{1}$, $C_{2}$) refer to coin flips, edges give the probabilities of moving to the corresponding node.
@@
In order to simulate $p_{1}$- and $p_{2}$-coins in the broadest set of Markov switching models, the skeleton $\breve{z}_{(\dot{\tau},\ddot{\tau})}$ of the bridge $z_{(\dot{\tau},\ddot{\tau})}$ must satisfy three requirements. Firstly, it must incorporate all previously evaluated locations on the path, such that all retrospective simulations are mutually consistent. Secondly, it must provide lower and upper bounds
$\infty<{z}^{\downarrow}_{(\dot{\tau},\ddot{\tau})}\leq z_{t}\leq{z}^{\uparrow}_{(\dot{\tau},\ddot{\tau})}<\infty,\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (19)
@@
which can then be propagated to bounds on $\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\varphi_{\theta}(\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),y_{\dot{\tau}})$. These are necessary to bound the time integral $\int_{\dot{\tau}}^{\ddot{\tau}}\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\mathop{}\!\mathrm{d}t$ and thereby the posterior odds, which is a prerequisite to constructing the factorization in (17). An upper bound on the time integral is also necessary for unbiased estimation of the intractable coin probabilities. Thirdly, it must be possible to recursively simulate $z_{(\dot{\tau},\ddot{\tau})}$ at an additional finite set of times according to the initial proposal distribution, conditional on the skeleton $\breve{z}_{(\dot{\tau},\ddot{\tau})}$. In this paper, the proposal distribution will always be the dominating measure $\mathbb{B}_{(\dot{\tau},\ddot{\tau})}$. Modulo these requirements, any method may in principle be used to generate the coins. The most general-purpose representations we’re aware of are the Bessel layer representation by [5] and the $\epsilon$-strong layer representation by [39]. Both representations rely on restricting the state space of $z$ to a random compact subset of $\mathbf{R}$, with the latter being a richer representation, but the former being more numerically reliable and easier to implement. Therefore, we use the Bessel layer representation in practice, as seen for example in [22].
In what follows, we merely presume that some lower and upper bounds ${z}^{\downarrow}_{(\dot{\tau},\ddot{\tau})}$ and ${z}^{\uparrow}_{(\dot{\tau},\ddot{\tau})}$ are available. Defining
$\mathcal{X}_{(\dot{\tau},\ddot{\tau})}=[\rho_{\theta}(y_{\dot{\tau}}){z}^{\downarrow}_{(\dot{\tau},\ddot{\tau})}+\eta_{\theta}(v_{\dot{\tau}})\land\eta_{\theta}(v_{\ddot{\tau}}),\rho_{\theta}(y_{\dot{\tau}}){z}^{\uparrow}_{(\dot{\tau},\ddot{\tau})}+\eta_{\theta}(v_{\dot{\tau}})\lor\eta_{\theta}(v_{\ddot{\tau}})]$ (20)
and observing that $x_{t}\in\mathcal{X}_{(\dot{\tau},\ddot{\tau})}$, we may define
@@
${\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\inf_{a\in\mathcal{X}_{(\dot{\tau},\ddot{\tau})}}\varphi_{\theta}(a,y_{\dot{\tau}}),\quad{\tilde{\varphi}}^{\uparrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\sup_{a\in\mathcal{X}_{(\dot{\tau},\ddot{\tau})}}\varphi_{\theta}(a,y_{\dot{\tau}}),$ (21)
@@
such that ${\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\leq\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\leq{\tilde{\varphi}}^{\uparrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$. Solving for the infimum and supremum (or bounds thereon) is a model-specific task, but may be done automatically by symbolic algebra packages.
@@ 3.2 Hidden Data Update
where $Z_{(\dot{\tau},\ddot{\tau})}^{\dagger}\sim\mathbb{B}_{(\dot{\tau},\ddot{\tau})}$ and $\kappa(z_{(\dot{\tau},\ddot{\tau})}^{\dagger})=1$, with respect to $\mathbb{B}_{(\dot{\tau},\ddot{\tau})}$. Using a Brownian bridge proposal ensures that the path skeletons can be simulated and extended, e.g. by the EA algorithm. $Y^{\dagger}$ is proposed independently from its prior distribution, i.e. $\kappa(y^{\dagger})=\pi(y^{\dagger}|\lambda)$. This is simply the forward measure of the jump process, and easily simulated from as in [26]. Given $Y^{\dagger}$, the proposal for $V_{r^{\dagger}}^{\dagger}$ is most readily understood in terms of $X_{r^{\dagger}}^{\dagger}=\eta_{\theta}(V_{r^{\dagger}}^{\dagger})$. We propose $X_{r^{\dagger}}^{\dagger}$ according to the dominating SDE $\mathop{}\!\mathrm{d}X_{t}=\rho_{\theta}(Y_{t}^{\dagger})\mathop{}\!\mathrm{d}W_{t}$ with induced conditional measure $\mathbb{M}|(x_{s},y^{\dagger},\theta)$. By the Markov property,
@@
[…]
@@
Thus, the proposal density on $V_{r^{\dagger}}^{\dagger}$ is given by $\kappa(v_{r^{\dagger}}^{\dagger}|v_{s},y^{\dagger})=\prod_{(\dot{s}\sim\ddot{s})\in s}\kappa(v_{r^{\dagger}\cap(\dot{s},\ddot{s})}^{\dagger}|v_{\{\dot{s},\ddot{s}\}},y_{[\dot{s},\ddot{s}]}^{\dagger})$. We spell out the step-by-step simulation routine below.
Algorithm 1 Algorithm for generating proposal from $\kappa(v_{r\cap(\dot{s},\ddot{s})}|v_{\{\dot{s},\ddot{s}\}},y_{[\dot{s},\ddot{s}]})$. $\mathbb{W}$ denotes the Wiener measure.
@@
$d_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\exp\left\{(\dot{\tau}-\ddot{\tau}){\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\right\},$ (27)
@@
$q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\exp\left\{\int_{\dot{\tau}}^{\ddot{\tau}}{\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\mathop{}\!\mathrm{d}t\right\}$ (28)
@@
such that $\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta^{\dagger})=(d_{\theta}\times q_{\theta})(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$, and where $q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\in[0,1]$. On that basis, the acceptance odds can be expressed as